The flywheel of a steam engine of mass 1000kg has radius of gyration 1m. If the maximum and minimum speed of the flywheel is 80r.p.m and 78r.m.p respectively. Find the fluctuation of energy


1
Expert's answer
2021-02-22T10:25:14-0500

"I=mr^2=1000\cdot1^2=1000\ (kg\cdot m^2)"


"80\ rpm=1.3333333\ s^{-1}"


"78\ rpm=1.3\ s^{-1}"


"E_1=I\omega_1^2/2=1000\cdot(2\cdot3.14\cdot1.3333333)^2/2=35056\ (J)"


"E_2=I\omega_2^2/2=1000\cdot(2\cdot3.14\cdot1.3)^2/2=33325\ (J)"


"\Delta E=35056-33325=1731\ (J)" . Answer